# TathyaNyaya and FactLegalLlama: Advancing Factual Judgment Prediction and Explanation in the Indian Legal Context

### v1

@@ 5.2 Training FactLegalLlama
The FactLegalLlama model, based on the LLaMa-3-8B architecture, was fine-tuned specifically for the FJPE task using NyayaFacts. The training process involved instruction-tuning with a diverse set of 16 templates designed to guide the model in judgment prediction and explanation tasks. We utilized low-rank adaptation (LoRA) to optimize model training on limited computational resources. Training parameters, such as quantization to 4-bit precision and gradient accumulation, ensured efficient usage of resources while maintaining model performance.
To further enhance its capabilities, FactLegalLlama was fine-tuned with both prediction-only and prediction-with-explanation tasks, enabling it to handle a wide range of factual judgment scenarios. The fine-tuning process emphasized the use of simplified prompts to ensure clarity and relevance in the generated outputs.
@@ 5.3.1 Language Model-Based Approach
For baseline comparisons, we utilized transformer-based models like InLegalBERT Paul et al. (2023), and XLNet Large Yang et al. (2019) for binary classification. Due to the token length constraints of these models, we adopted a chunking strategy by dividing documents into 512-token segments with a 100-token overlap to preserve context. Chunk-level predictions were aggregated to generate final case-level predictions.
@@ 5.3.2 Large Language Model-based Approach
We utilized FactLegalLlama, our instruction-tuned LLaMa-3-8B model Dubey et al. (2024), for judgment prediction-only instructions, where the model predicts judicial outcomes solely based on the factual inputs. The training data from TathyaNyaya was used to train the factual prediction context, emphasizing precision.
5.4 Fact-Based Judgment Prediction with Explanation (FJPE)
@@ 5.5 Prompts Used
Prompts for both prediction and explanation tasks were carefully designed and adapted from previous studies Vats et al. (2023); Nigam et al. (2024c, d). For prediction-only tasks, the prompts instructed the model to output a binary decision. For prediction-with-explanation tasks, the prompts included directives to explain the reasoning behind the prediction. These templates are detailed in Table 6 in the Appendix.
@@ 6 Evaluation Metrics
To rigorously assess the performance of our models on judgment prediction and factual explanations in the TathyaNyaya test dataset, we employed a suite of evaluation metrics. For judgment prediction, we report Macro Precision, Macro Recall, Macro F1, and Accuracy. For evaluating the quality of explanations, both quantitative and qualitative methods were applied.
@@
    Lexical-Based Evaluation: We used traditional lexical similarity metrics, including ROUGE (ROUGE-1, ROUGE-2, and ROUGE-L) Lin (2004), BLEU Papineni et al. (2002), and METEOR Banerjee and Lavie (2005). These metrics measure word overlap and sequence alignment between generated explanations and reference texts, providing a quantitative measure of the accuracy of lexical content.
@@
    Semantic Similarity Evaluation: To assess the semantic alignment of the generated explanations, we applied BERTScore Zhang et al. (2020), which evaluates semantic similarity between the generated text and reference explanations. Additionally, BLANC Vasilyev et al. (2020) was utilized to estimate the contextual relevance and coherence of the generated text in the absence of a gold-standard reference.
@@
    Inter-Annotator Agreement: In the construction of NyayaFacts, each document was annotated by a single annotator due to the complexity and scale of the dataset. While this approach ensured a manageable workflow and timely completion, it precludes the direct calculation of inter-annotator agreement. As a result, we do not report inter-annotator metrics, and future work may consider sampling subsets of the data for multiple annotations to facilitate such evaluations.
@@ 7 Results and Analysis
FactLegalLlama 0.5416 0.5312 0.5036 0.5386 NyayaFacts Single
@@
FactLegalLlama 0.5390 0.5368 0.5318 0.5401 NyayaFacts Multi
@@
FactLegalLlama 0.5139 0.5122 0.4922 0.5042 NyayaScrape Single
@@
FactLegalLlama 0.4951 0.4966 0.4516 0.4884 NyayaScrape Multi
@@
FactLegalLlama 0.5574 0.5372 0.5191 0.6045 NyayaScrape Single
@@
FactLegalLlama 0.5439 0.5317 0.5177 0.5877 NyayaScrape Multi
@@
FactLegalLlama 0.5832 0.5868 0.5792 0.5840 NyayaFacts Single
@@
FactLegalLlama 0.6541 0.6583 0.6552 0.6651 NyayaFacts Multi
Table 3: Performance metrics of models evaluated on NyayaFacts and NyayaScrape test data. Each block shows results obtained by training on either NyayaFacts or NyayaScrape data (single or multi variants), then testing on corresponding subsets. The best scores in each section are highlighted in bold.
@@
InLegalBert 0.5870 0.5857 0.5864 0.5885 NyayaFacts Single
@@
InLegalBert 0.5886 0.5560 0.5719 0.5421 NyayaFacts Multi
@@
InLegalBert 0.5342 0.5180 0.5260 0.5023 NyayaScrape Single
@@
InLegalBert 0.5789 0.5409 0.5592 0.5249 NyayaScrape Multi
@@
InLegalBert 0.6199 0.6197 0.6198 0.6167 NyayaFacts Single
@@
InLegalBert 0.6222 0.5986 0.6102 0.5839 NyayaFacts Multi
@@
InLegalBert 0.5760 0.5311 0.5526 0.5061 NyayaScrape Single
@@
InLegalBert 0.5659 0.5215 0.5428 0.4950 NyayaScrape Multi
@@
Table 4: Model performance on NyayaFilter and NyayaSimplify test datasets. For NyayaFilter, results illustrate how automatically retrieved factual data affects performance when models are trained on NyayaFacts or NyayaScrape datasets. For NyayaSimplify, results show the impact of paraphrasing complex legal texts into simpler language. Bolded scores indicate the best performance in each section.
@@
Training Data Testing Data Rouge-1 Rouge-2 Rouge-L BLEU METEOR
BERT Score
BLANC
No Training NyayaFacts 0.2757 0.0998 0.1448 0.0395 0.1650 0.5250 0.0814
No Training NyayaScrape 0.1877 0.0750 0.1297 0.0371 0.1848 0.4819 0.0927
NyayaFacts Single NyayaFacts 0.3216 0.1085 0.1897 0.0419 0.1798 0.5785 0.0958
NyayaFacts Multi NyayaFacts 0.3383 0.1133 0.1950 0.0483 0.2120 0.5843 0.1031
NyayaScrape Single NyayaScrape 0.1172 0.0480 0.0864 0.0211 0.0967 0.3927 0.0609
NyayaScrape Multi NyayaScrape 0.1720 0.0762 0.1269 0.0314 0.1287 0.4516 0.0782
NyayaSimplify NyayaSimplify 0.2818 0.0764 0.1816 0.0242 0.1689 0.5559 0.0731
Table 5: Performance of FactLegalLlama on the FJPE task. The base model is LLaMa-3-8B. "No Training" indicates results from the unmodified (vanilla) model. Other rows show improvements after fine-tuning with different subsets of the TathyaNyaya data. Bolded values represent the best performance within a given evaluation scenario.
In this section, we present and interpret the performance of our models across various datasets and experimental settings. We focus first on raw judgment prediction results using NyayaFacts and NyayaScrape data, then on the performance improvements or trade-offs observed in the NyayaFilter and NyayaSimplify settings. Finally, we analyze the explanation quality generated by FactLegalLlama using both lexical and semantic metrics.
@@ 7.1 Performance on NyayaFacts and NyayaScrape
We begin by examining model performances on the NyayaFacts and NyayaScrape test sets, as reported in Table 3. Each model (InLegalBERT, XLNet_Large, and FactLegalLlama) was evaluated under different training configurations, including Single and Multi.
@@
Across both NyayaFacts and NyayaScrape test sets, XLNet_Large consistently outperforms InLegalBERT on macro Precision, Recall, F1, and Accuracy metrics. For instance, when trained on NyayaFacts Single, XLNet_Large achieves a macro F1 of 0.6052 and Accuracy of 0.6061, surpassing InLegalBERT’s macro F1 of 0.5935 and Accuracy of 0.5932. This trend persists in most training and testing configurations, highlighting XLNet_Large’s robust capability for factual judgment prediction in the given domain.
@@
FactLegalLlama, while instruction-tuned for outcome prediction, lags behind the transformer-based baselines in raw prediction performance. For example, when trained on NyayaFacts Single and tested on NyayaFacts, it obtains a macro F1 of 0.5036 compared to XLNet_Large’s 0.6052. A similar gap is observed across other splits. Although FactLegalLlama underperforms in direct classification metrics, its strength lies in generating explanations, as discussed later.
@@ 7.2 Impact of Fact Retrieval (NyayaFilter) and Text Simplification (NyayaSimplify)
When comparing NyayaFilter results to the original NyayaFacts and NyayaScrape sets, we see that while performance can fluctuate, some models benefit from training on data where fact and non-fact segments are clearly distinguished. For example, on the NyayaFilter test set derived from NyayaFacts Single, InLegalBERT attains a macro F1 of 0.5864, maintaining competitive performance. XLNet_Large, although not always the top performer here, still sustains a strong baseline. These findings suggest that automatically retrieved factual subsets can be used without severely degrading model performance.
@@
Paraphrasing complex legal language into simpler text (the NyayaSimplify scenario) generally helps models retain or slightly improve performance. For instance, with NyayaFacts Single, InLegalBERT reaches a macro F1 of 0.6198 and XLNet_Large hits an Accuracy of 0.6200 on the simplified data, both representing small yet noteworthy improvements compared to their performance on the original complex texts. This trend indicates that reducing linguistic complexity can aid models in understanding and classifying factual statements more accurately.
7.3 Quality of Explanations from FactLegalLlama
Table 5 presents the evaluation of FactLegalLlama on the explanation generation task, measured through both lexical (Rouge, BLEU, METEOR) and semantic (BERTScore, BLANC) metrics. We compare a "No Training" scenario (using the LLaMa-3-8B model) with fine-tuned versions on different subsets of TathyaNyaya data.
@@ 7.3 Quality of Explanations from FactLegalLlama
Fine-tuning LLaMa-3-8B (FactLegalLlama) on factual data substantially improves its explanation quality. For NyayaFacts, training on the Multi subset yields the strongest results, with Rouge-1 at 0.3383 and a BERTScore of 0.5843, outperforming both the "No Training" scenario and the Single subset training. This suggests that exposure to more complex, multi-petition cases helps the model generate richer, more contextually sensitive explanations.
@@ 8 Conclusions and Future Work
We introduced TathyaNyaya, a fact-focused dataset for judgment prediction and explanation within the Indian legal domain, and FactLegalLlama, an instruction-tuned model delivering fact-grounded rationales. By emphasizing factual content rather than full judgments, TathyaNyaya aligns more closely with actual legal decision-making scenarios, while FactLegalLlama highlights the value of coupling predictive accuracy with transparent explanations. Preprocessing steps such as fact filtering and paraphrasing further enhance model clarity and performance, and domain-specific fine-tuning proves essential for capturing legal subtleties. Future work may extend these findings to other jurisdictions, refine fact extraction techniques, integrate, and interpretability frameworks. These efforts collectively advance transparent, accessible, and reliable AI-assisted judicial processes.

### v2

@@ 5.2 Training FactLegalLlama
The FactLegalLlama model, based on LLaMa-3-8B architecture, was fine-tuned specifically for the FJPE task using NyayaFacts. The training process involved instruction-tuning with a diverse set of 16 templates designed to guide the model in judgment prediction and explanation tasks. We utilized low-rank adaptation (LoRA) to optimize model training on limited computational resources. Training parameters, such as quantization to 4-bit precision and gradient accumulation, ensured efficient usage of resources while maintaining model performance. To further enhance its capabilities, FactLegalLlama was fine-tuned with both prediction-only and with explanation tasks, enabling it to handle a wide range of factual judgment scenarios.
@@ 5.3.1 Language Model-Based Approach
For baseline comparisons, we utilized transformer-based models like InLegalBERT paul-2022-pretraining, and XLNet Large yang2019xlnet for binary classification. Due to the token length constraints of these models, we adopted a chunking strategy by dividing documents into 512-token segments with a 100-token overlap to preserve context. Chunk-level predictions were aggregated to generate final case-level predictions.
@@ 5.3.2 Large Language Model-based Approach
We utilized FactLegalLlama, our instruction-tuned LLaMa-3-8B model dubey2024llama, for judgment prediction-only instructions, where the model predicts judicial outcomes solely based on the factual inputs. The training data from TathyaNyaya was used to train the factual prediction context, emphasizing precision.
5.4 Prediction with Explanation (FJPE)
@@ 5.5 Prompts Used
Prompts for both prediction and explanation tasks were carefully designed the prompts. For prediction-only tasks, the prompts instructed the model to output a binary decision. For prediction-with-explanation tasks, the prompts included directives to explain the reasoning behind the prediction. Templates are detailed in Table 8 in the Appendix.
@@ 6 Evaluation Metrics
To rigorously assess the performance of our models on judgment prediction and factual explanations in the TathyaNyaya test dataset, we employed a suite of evaluation metrics. For judgment prediction, we report Macro Precision, Recall, F1, and Accuracy. For evaluating the quality of explanations, both quantitative and qualitative methods were applied.
@@
    Lexical-Based Evaluation: We used traditional lexical similarity metrics, including ROUGE-1/2/L lin-2004-rouge, and BLEU papineni-etal-2002-bleu. These metrics measure word overlap and sequence alignment between generated explanations and reference texts, providing a quantitative measure of the accuracy of lexical content.
@@
    Semantic Similarity Evaluation: To assess the semantic alignment of the generated explanations, we applied BERTScore BERTScore, which evaluates semantic similarity between the generated text and reference explanations. Additionally, BLANC blanc was utilized to estimate the contextual relevance and coherence of the generated text in the absence of a gold-standard reference.
@@
    Expert Evaluation: To validate the interpretability and legal soundness of the model-generated explanations, we conduct an expert evaluation involving legal professionals. They rate a representative subset of the generated outputs on a 1–10 Likert scale across three criteria: factual accuracy, legal relevance, and completeness of reasoning. A score of 1 denotes a poor or misleading explanation, while a 10 reflects high legal fidelity and argumentative soundness. This evaluation provides critical insights beyond automated metrics.
  4. 4.
    Inter-Annotator Agreement (IAA): To further ensure the reliability and consistency of expert judgments, we computed multiple standard inter-rater agreement metrics. Specifically, Fleiss’ Kappa fleiss1971measuring was used to evaluate the overall agreement among multiple raters, Cohen’s Kappa cohen1960coefficient captured pairwise agreement while adjusting for chance, and the Intraclass Correlation Coefficient (ICC) shrout1979intraclass measured the reliability of continuous ratings across raters. In addition, Krippendorff’s Alpha krippendorff2018content provided a robust measure suitable for ordinal scales and missing data, while the Pearson Correlation Coefficient benesty2009pearson quantified the linear consistency of expert scores. The results demonstrated substantial agreement across these different measures, reinforcing the credibility of the evaluation and lending strong support to the robustness of our findings.
@@ 7 Results and Analysis
FactLegalLlama 0.5416 0.5312 0.5036 0.5386
NyayaFacts
Single
@@
FactLegalLlama 0.5390 0.5368 0.5318 0.5401
NyayaFacts
Multi
@@
FactLegalLlama 0.5139 0.5122 0.4922 0.5042
NyayaScrape
Single
@@
FactLegalLlama 0.4951 0.4966 0.4516 0.4884
NyayaScrape
Multi
@@
FactLegalLlama 0.5574 0.5372 0.5191 0.6045
NyayaScrape
Single
@@
FactLegalLlama 0.5439 0.5317 0.5177 0.5877
NyayaScrape
Multi
@@
FactLegalLlama 0.5832 0.5868 0.5792 0.5840
NyayaFacts
Single
@@
FactLegalLlama 0.6541 0.6583 0.6552 0.6651
NyayaFacts
Multi
Table 4: Performance metrics of models evaluated on NyayaFacts and NyayaScrape test data. Each block shows results obtained by training on either NyayaFacts or NyayaScrape data (single or multi variants), then testing on corresponding subsets.
@@
InLegalBert 0.5870 0.5857 0.5864 0.5885
NyayaFacts
Single
@@
InLegalBert 0.5886 0.5560 0.5719 0.5421
NyayaFacts
Multi
@@
InLegalBert 0.5342 0.5180 0.5260 0.5023
NyayaScrape
Single
@@
InLegalBert 0.5789 0.5409 0.5592 0.5249
NyayaScrape
Multi
@@
InLegalBert 0.6199 0.6197 0.6198 0.6167
NyayaFacts
Single
@@
InLegalBert 0.6222 0.5986 0.6102 0.5839
NyayaFacts
Multi
@@
InLegalBert 0.5760 0.5311 0.5526 0.5061
NyayaScrape
Single
@@
InLegalBert 0.5659 0.5215 0.5428 0.4950
NyayaScrape
Multi
@@
Table 5: Model performance on NyayaFilter and NyayaSimplify test datasets. For NyayaFilter, results illustrate how automatically retrieved factual data affects performance when models are trained on NyayaFacts or NyayaScrape datasets. For NyayaSimplify, results show the impact of paraphrasing complex legal texts into simpler language.
@@
Training
Data
Testing
Data
R1 R2 RL BLEU BERTScore BLANC
Expert
Score
No Training NyayaFacts 0.28 0.10 0.14 0.04 0.53 0.08 4.2
No Training NyayaScrape 0.19 0.08 0.13 0.04 0.48 0.09 4.0
NyayaFacts Single NyayaFacts 0.32 0.11 0.19 0.04 0.58 0.10 7.6
NyayaFacts Multi NyayaFacts 0.34 0.11 0.20 0.05 0.58 0.10 8.3
NyayaScrape Single NyayaScrape 0.12 0.05 0.09 0.02 0.39 0.06 4.9
NyayaScrape Multi NyayaScrape 0.17 0.08 0.13 0.03 0.45 0.08 5.6
NyayaSimplify NyayaSimplify 0.28 0.08 0.18 0.02 0.56 0.07 6.7
Table 6: Performance of FactLegalLlama on the FJPE task. The base model is LLaMa-3-8B. "No Training" indicates results from the unmodified (vanilla) model. Other rows show improvements after fine-tuning with different subsets of the TathyaNyaya data.
In this section, we present and interpret the performance of our models across various datasets and experimental settings. We focus first on raw judgment prediction results using NyayaFacts and NyayaScrape data, then on the performance improvements or trade-offs observed in the NyayaFilter and NyayaSimplify settings. Finally, we analyze the explanation quality generated by using lexical, semantic, and expert evaluation metrics.
@@ 7.1 Performance on NyayaFacts and NyayaScrape
We begin by examining model performances on the NyayaFacts and NyayaScrape test sets, as reported in Table 4. Each model was evaluated under different training configurations, including Single and Multi.
@@
Across both NyayaFacts and NyayaScrape test sets, XLNet large consistently outperforms InLegalBERT. For instance, when trained on NyayaFacts Single, XLNet achieves a macro F1 of 0.6052 and Accuracy of 0.6061, surpassing InLegalBERT’s macro F1 of 0.5935 and Accuracy of 0.5932. This trend persists in most training and testing configurations, highlighting XLNet’s robust capability for factual judgment prediction in the given domain.
@@
FactLegalLlama, while instruction-tuned for outcome prediction, lags behind the transformer-based baselines in raw prediction performance. For example, when trained on NyayaFacts Single and tested on NyayaFacts, it obtains a macro F1 of 0.5036 compared to XLNet_Large’s 0.6052. A similar gap is observed across other splits.
@@ 7.2 Impact of Fact Retrieval (NyayaFilter) and Text Simplification (NyayaSimplify)
When comparing NyayaFilter results to the original NyayaFacts and NyayaScrape sets, we see that while performance can fluctuate, some models benefit from training on data where fact and non-fact segments are clearly distinguished. For example, on the NyayaFilter test set derived from NyayaFacts Single, InLegalBERT attains a macro F1 of 0.5864, maintaining competitive performance. These findings suggest that automatically retrieved factual subsets can be used without severely degrading model performance.
@@
Paraphrasing complex legal language into simpler text (the NyayaSimplify scenario) generally helps models retain or slightly improve performance. For instance, with NyayaFacts Single, InLegalBERT reaches a macro F1 of 0.6198 and XLNet_Large hits an Accuracy of 0.6200 on the simplified data, both representing small yet noteworthy improvements compared to their performance on the original complex texts. This trend indicates that reducing linguistic complexity can aid models in understanding and classifying factual statements better.
Base Model Training Data Testing Data Fleiss’ $\kappa$ Cohen’s $\kappa$ ICC Kripp. $\alpha$ Pearson Corr.
Meta-LLaMA-3-8B No Training NyayaFacts 0.62 0.59 0.64 0.61 0.65
Meta-LLaMA-3-8B No Training NyayaScrape 0.54 0.50 0.55 0.53 0.56
Meta-LLaMA-3-8B NyayaFacts Single NyayaFacts 0.70 0.68 0.71 0.69 0.72
Meta-LLaMA-3-8B NyayaFacts Multi NyayaFacts 0.81 0.79 0.83 0.80 0.84
Meta-LLaMA-3-8B NyayaScrape Single NyayaScrape 0.58 0.54 0.60 0.57 0.59
Meta-LLaMA-3-8B NyayaScrape Multi NyayaScrape 0.63 0.60 0.66 0.62 0.64
Meta-LLaMA-3-8B NyayaSimplify NyayaSimplify 0.69 0.66 0.70 0.68 0.70
Table 7: Inter-Annotator Agreement (IAA) metrics for expert evaluation across different training and testing setups. Higher scores indicate stronger agreement and reliability of expert-based assessments.
7.3 Quality of Explanations
Table 6 presents the evaluation of FactLegalLlama on the explanation generation task, measured through lexical, semantic, and expert evaluation metrics. We compare a "No Training" scenario with fine-tuned versions of TathyaNyaya data.
@@ 7.3 Quality of Explanations
FactLegalLlama on factual data substantially improves its explanation quality. For NyayaFacts, training on the Multi subset yields the strongest results, outperforming both the "No Training" scenario and the Single subset training. This suggests that exposure to more complex, multi-petition cases helps the model generate richer, more contextually sensitive explanations.
@@
Expert Evaluation:
As shown in Table 6, we complemented our automatic lexical and semantic evaluation with a human-centric expert assessment using a 1–10 Likert scale. Legal experts rated the generated explanations on clarity, legal relevance, and factual consistency. Results demonstrate that FactLegalLlama fine-tuned on NyayaFacts Multi, achieved the highest score (average 8.3), reflecting strong alignment with human-annotated rationales. Fine-tuning on simplified data (NyayaSimplify) and single-case subsets also yielded substantial improvements (6.7 and 7.6, respectively), while explanations from the zero-shot model (No Training) were rated notably lower (around 4.0). These findings highlight the importance of domain-specific supervision and exposure to complex multi-petition cases for producing high-quality, interpretable explanations.
To ensure reliability of expert ratings, we further computed IAA results, summarized in Table 7, which indicate moderate to high agreement across evaluators, with the strongest consistency observed for NyayaFacts Multi. This demonstrates that expert judgments were stable and reproducible, reducing subjectivity and strengthening the credibility of our evaluation framework. Together, the expert scores and IAA results confirm that improvements in explanation quality are not only measurable but also consistently recognized by multiple evaluators.
@@ 8 Conclusions and Future Work
We introduced TathyaNyaya, a fact-focused dataset for judgment prediction and explanation within the Indian legal domain, and FactLegalLlama, an instruction-tuned model delivering fact-grounded rationales. By emphasizing factual content rather than full judgments, TathyaNyaya aligns more closely with actual legal decision-making scenarios, while FactLegalLlama highlights the value of coupling predictive accuracy with transparent explanations. Preprocessing steps such as fact filtering and paraphrasing further enhance model clarity and performance, and domain-specific fine-tuning proves essential for capturing legal subtleties. Future work may extend these findings to other jurisdictions.
Acknowledgements
We would like to express our sincere gratitude to the anonymous reviewers for their constructive feedback and valuable suggestions, which greatly improved the quality of this paper. We also thank the student research assistants and legal experts from various law colleges for their dedicated efforts in annotation and expert evaluation. Their contributions were instrumental in ensuring the quality and reliability of the dataset and evaluation process. We gratefully acknowledge the support of BharatGen, India, for providing access to computational resources and hardware infrastructure used in this research. Their assistance played a vital role in model training and large-scale experimentation. The majority of this work was conducted while the first author was affiliated with the Indian Institute of Technology Kanpur. The author is currently affiliated with the University of Birmingham Dubai.